# Precise Attribute Intensity Control in Large Language Models via Targeted Representation Editing

### v1

@@ Abstract
Precise attribute intensity control—generating Large Language Model (LLM) outputs with specific, user-defined attribute intensities—is crucial for AI systems adaptable to diverse user expectations. Current LLM alignment methods, however, typically provide only directional or open-ended guidance, failing to reliably achieve exact attribute intensities. We address this limitation with three key designs: (1) reformulating precise attribute intensity control as a target-reaching problem, rather than simple maximization; (2) training a lightweight value function via temporal-difference learning to predict final attribute intensity scores from partial generations, thereby steering LLM outputs; and (3) employing gradient-based interventions on hidden representations to navigate the model precisely towards specific attribute intensity targets. Our method enables fine-grained, continuous control over attribute intensities, moving beyond simple directional alignment. Experiments on LLaMA-3.2-3b and Phi-4-mini confirm our method’s ability to steer text generation to user-specified attribute intensities with high accuracy. Finally, we demonstrate efficiency enhancements across three downstream tasks: preference data synthesis, Pareto frontier approximation and optimization, and distillation of aligned behaviors for intervention-free inference. Our code is available on https://github.com/Pre-Control/pre-control.
@@ 1 Introduction
Precise control over attribute intensity is critical for tailoring large language model (LLM) outputs to diverse contexts and user needs [2, 26]. Rather than merely pushing attributes in a single direction, precise attribute intensity control enables fine-grained adjustment of text attributes—such as tone, helpfulness, or formality—on a continuous scale [7, 28]. This capability is essential for practical applications, such as calibrating an email’s tone from slightly formal for a colleague to highly formal for an executive [6]. The stakes are even higher in multi-objective alignment, where attributes conflict with each other [4, 43]. Navigating trade-offs between attributes, such as maximizing helpfulness while minimizing misinformation, requires scalar-level adjustments to identify optimal compromises [2, 40]. However, adjusting an LLM along continuous attribute trade-offs is difficult. While sophisticated prompting can elicit complex behaviors, it remains an unreliable method for precise and reproducible attribute control. The mapping from a qualitative description to a point in the model’s attribute space is non-trivial and highly sensitive to phrasing. This indirect mechanism makes it challenging to achieve specific scalar targets of the attribute intensity, especially when attributes are entangled in multi-objective scenarios [10, 30].
Existing alignment paradigms fundamentally lack the capability for efficient precise attribute intensity control. Fine-tuning methods like Reinforcement Learning from Human Feedback (RLHF; [30, 46, 33]) and direct preference optimization (DPO; [28]) produce static models that capture an average of desired behaviors, requiring expensive retraining to shift priorities. While recent advances in multi-objective alignment [29, 45, 41, 44] can identify Pareto-optimal solutions, they often require extensive training to approximate a global Pareto set rather than enabling efficient, controllable projection of generations onto specific points on that frontier. Test-time methods avoid retraining but have their own limitations. Prompting approaches [1, 42, 18] rely entirely on the model’s interpretation of style instructions, yielding inconsistent results. Guided decoding [14, 23, 12, 13] typically treat attribute intensity as categorical rather than continuous. Moreover, without modifying the model’s parameters, these methods remain constrained by the pretrained model’s capabilities, making effective fine-grained control (e.g., adjusting helpfulness=4=4, complexity=2=2 on a 040-4 scale) unattainable.
@@ 2.1 From Standard LLM Alignment To Target Reaching Formulation
$\vskip 2.15277pt\max_{\theta}\mathbb{E}_{x\sim\pi_{\theta}}[R(x)],$ (1)
@@ 3.1 Value Function Training via Temporal Difference Learning
In this formulation, $s_{t+n}$ denotes the state reached $n$ steps after $s_{t}$, where $T$ represents the total sequence length. The term $V_{\phi}(s_{t+n})$ serves as a bootstrap estimate of the eventual terminal score starting from that future state, while $r^{T}$ is the final, episode-level reward for the completed sequence. The decay factor $\lambda$ trades off short-horizon bootstrapping against reliance on the terminal Monte Carlo target—approaching pure MC as $\lambda\to 1$. The value function is then trained to minimize the mean squared error between its predictions and the generalized returns:
@@ 3.2 Test-time Intervention for Target Attribute Intensity Control
With a trained value function that can predict attribute intensity scores from partial generations, we leverage it to guide the policy model toward generating text with a specific attribute score through targeted representation editing. Unlike previous approaches [13, 16, 18, 42] that merely push the model to maximize or minimize a preference, Pre-Control enables precise targeting of any scores within the full range of attribute intensities.
@@ 3.3 Efficient Pareto Frontier Approximation
$\mathbf{t}=I(\mathbf{n}_{1},\mathbf{n}_{2},\beta),\quad\beta\in[0,1]$ (11)
@@
The core of our approach is using our precise attribute intensity control capability to directly generate samples at specific target points along the Pareto frontier, which traditional methods cannot achieve.
@@
For each iteration, we identify the most promising target by calculating the coverage gap at each point $\mathbf{t}$ as:
$G(\mathbf{t},\mathcal{N})=\min_{\mathbf{n}\in\mathcal{N}}|\mathbf{t}-\mathbf{n}|_{2}.$ (12)
@@ 4.2 Main Results
We evaluate the effectiveness of Pre-Control for precise attribute intensity control on HelpSteer2 and Code-UltraFeedback. Table 1 presents the main results on both relative positive and negative target vectors, which illustrate Pre-Control’s bidirectional finer-grained control capability. Crucially, the strong performance of our method is not limited to these specific points. We provide a comprehensive evaluation across a wide range of target scores in Appendix B.3, with full results in Table 11 and  12, which confirms the robustness and consistency of our findings. To better demonstrate the effectiveness of our approach, we present a comparison of attribute intensity distributions before and after our intervention in Figure 5.
@@ 5.1 LLM Alignment
Alignment approaches for LLMs fall into two primary paradigms. Fine‐tuning via RLHF [30, 46, 33]—where a reward model guides policy optimization—yields robust performance but depends on a multi‐stage loop of reward learning, policy updates, and rollouts, which can be resource‐intensive [30, 33]. Direct Preference Optimization (DPO) [28] recasts this as a supervised loss, removing the need for online sampling, yet still demands significant memory to maintain both policy and reference models. Inference‐time interventions sidestep model updates: prompt engineering—crafting instructions (with or without examples)—can nudge outputs toward desired behaviors with almost no extra compute [1]. Guided decoding as another effective branch has also been well-explored: ARGS weave reward‐model scores into token probabilities to steer generation [14]. Mudgal et al.[23] and Han et al.[12] train a prefix‑based reward scorer to guide generation from a partial hypothesis. DeAL [13], by contrast, casts decoding as an A* search, using heuristic costs to optimize token selection. TransferQ [5] introduces an inference‐time policy adjustment that minimally shifts a pretrained model’s token distribution toward higher‐reward outputs via a quantile‐based optimization. However, these methods fundamentally lack a principled mechanism for precise attribute intensity control, as they are designed to monotonically shift outputs toward human-preferred extremes rather than targeting specific points along the preference spectrum.
@@
Another critical direction in LLM alignment is multi-objective alignment, which is crucial for real-world deployment where LLMs must balance competing attributes based on user preferences. Recent works on multi-objective alignment have explored various ways to approximate Pareto-optimal trade-offs. [29] trains separate policies for each reward preference via RLHF and interpolates them post hoc. MODPO [45] extends Direct Preference Optimization to handle multiple objectives. RiC [41] reduces training costs by applying reward-conditioned supervised fine-tuning and lightweight online data augmentation. Panacea [44] further embeds preference vectors into model parameters through SVD-LoRA, enabling a single model to generalize across objectives after training. Despite these advancements, these methods rely on costly retraining to inject various multi-objective preferences. In contrast, our method achieves efficient alignment entirely at test time.
@@ 5.2 Representation Engineering
Activation perturbation began with plug‑and‑play methods that use attribute‑specific classifiers to nudge hidden states toward desired traits [6]. Subsequent studies showed that both learned and handcrafted steering vectors can control style [31, 34] and that targeting attention‑head outputs boosts factual accuracy [17]. [27] applies steering vector, constructed from residual-stream activation differences between positive and negative exemplars, at inference to intervene behaviors. [3] optimizes steering vector using contrastive human-preference pair and use it to inject personalized control without additional model training. Liu et al. further interpret in‑context learning as shifting latent states toward task‑relevant regions [20]. More recently, representation fine‑tuning leverages low‑rank projection matrices to edit activations efficiently—often matching or outperforming parameter‑efficient tuning [39, 38]—whereas Liu et al.’s two‑phase approach first identifies steering directions via fine‑tuning before applying them, adding extra complexity [21]. Similarly, these methods primarily focus on binary or categorical attribute control rather than precisely targeting specific attribute intensities on a continuous scale.
@@ 7 Reproducibility Statement
We release code, configs, and scripts at https://github.com/Pre-Control/pre-control. The core algorithmic details are specified in Section 3 and Algorithm 1, and the full experimental setup appears in Section 4. Dataset descriptions and preprocessing steps for HelpSteer2 and Code-UltraFeedback are provided in Appendix C.1 and Appendix C.2, respectively. Implementation specifics—including model choices, intervention layer, value-function architecture, and training targets (ArmoRM), hyperparameters, random seeds, and decoding settings—are documented in Appendix C.3 and in the released configuration files. Our Pareto-frontier construction and interpolation procedures, along with the hypervolume, sparsity, and #PF computations, are detailed in Appendix C.4; metric definitions (Self-BLEU, $\ell_{1}$ distance to target, and Success Rate with filtering rules) are summarized in the Metrics subsection of Section 4 and mirrored in the repository’s evaluation scripts. Computing infrastructure (hardware, GPU hours, and environment) is reported in Appendix D.
@@ B.3 Full Intervention Results
To demonstrate the robustness of our approach, we assess its performance across a range of target attribute intensity scores. The complete results for these varied targets are shown in Table 11 and 12. Pre-Control demonstrates a consistent, strong performance compared to all baselines in various settings.
@@ C.4.2 Upper Convex Hull Interpolation
Appendix D Computing Infrastructure
@@
Appendix E Latency
To evaluate the runtime efficiency of Pre-Control, we record GPU usage for both value‐function training and inference‐time intervention on the hardware detailed in Section D. Training the value function on 1,0162 samples (with 519 held‑out test cases) requires only 0.34 GPU hours. Inference‑time interventions, whose cost varies with target attribute intensity and learning rate, incur an average of 0.47 GPU hours. Together, these measurements highlight the computational efficiency of Pre-Control.
Appendix F Case Study
In Table˜9 and Table˜10, we present qualitative examples demonstrating Pre-Control’s ability to precisely control attribute intensities.
@@ Appendix F Case Study
Positive Target Scenario. The base model generates a response with scores [4,4,4,1,1] for helpfulness, coherence, correctness, complexity, and verbosity, respectively. While the response is helpful and correct, it lacks detail—providing only minimal explanations without elaborating on command options or their purposes. To address this deficiency, we set target scores of [4,4,4,2,2], aiming to maintain the high quality while increasing both complexity and verbosity to provide more comprehensive information. After applying Pre-Control, the model successfully achieves these exact target scores by enriching the response with explicit clarifications of command flags, detailed option descriptions, and expanded explanations of each component’s purpose.
@@
Positive Target Example Prompt:

### v2

@@ Abstract
Precise attribute intensity control—generating Large Language Model (LLM) outputs with user-defined attribute intensities—is crucial for AI systems adaptable to diverse user expectations. Current LLM alignment methods, however, typically provide only directional or open-ended guidance, failing to achieve exact attribute intensities. We address this limitation with three key designs: (1) reformulating precise attribute intensity control as a target-reaching problem, rather than simple maximization; (2) training a lightweight value function via temporal-difference learning to predict final attribute intensity scores from partial generations, thereby steering LLM outputs; and (3) employing gradient-based interventions on hidden representations to navigate the model precisely towards specific attribute intensity targets. We enable fine-grained, continuous control over attribute intensities, moving beyond simple directional alignment. Experiments on LLaMA-3.2-3b and Phi-4-mini confirm our method’s ability to steer text generation to user-specified attribute intensities with high accuracy. Finally, we demonstrate efficiency enhancements across three downstream tasks: preference data synthesis, Pareto frontier approximation and optimization, and distillation of aligned behaviors for intervention-free inference. We release our code on https://github.com/Pre-Control/pre-control.
@@ 1 Introduction
Precise control over attribute intensity is critical for tailoring large language model (LLM) outputs to diverse contexts and user needs [3, 31]. Rather than merely pushing attributes in a single direction, precise attribute intensity control enables fine-grained adjustment of text attributes—such as tone, helpfulness, or formality—on a continuous scale [34]. This capability is essential for practical applications, such as calibrating an email’s tone from slightly formal for a colleague to highly formal for an executive [8]. The stakes are even higher in multi-objective alignment, where attributes conflict with each other [5, 49]. Navigating trade-offs between attributes, such as maximizing helpfulness while minimizing misinformation, requires scalar-level adjustments to identify optimal compromises [3, 46]. However, adjusting an LLM along continuous attribute trade-offs is difficult. While sophisticated prompting can elicit complex behaviors, it remains an unreliable method for precise and reproducible attribute control. The mapping from a qualitative description to a point in the model’s attribute space is non-trivial and highly sensitive to phrasing. This indirect mechanism makes it challenging to achieve specific scalar targets, especially when attributes are entangled in multi-objective scenarios [36, 11].
Existing alignment paradigms fundamentally lack the capability for efficient precise attribute intensity control. Fine-tuning methods like Reinforcement Learning from Human Feedback (RLHF; [36, 52, 39]) and direct preference optimization (DPO; [34]) produce static models that capture an average of desired behaviors, requiring expensive retraining to shift priorities. While recent advances in multi-objective alignment [35, 51, 47, 50] can identify Pareto-optimal solutions, they often require extensive training to approximate a global Pareto set rather than enabling efficient, controllable projection of individual generations onto specific points on that frontier. Test-time methods avoid retraining but have their own limitations. Prompting approaches [2, 48, 21] rely entirely on the model’s interpretation of style instructions, yielding inconsistent results. Guided decoding [16, 28, 14, 15] typically treat attribute intensity as categorical rather than continuous. Moreover, without modifying the model’s parameters, these methods remain constrained by the pretrained model’s capabilities, making effective fine-grained control (e.g., adjusting helpfulness=4=4, complexity=2=2 on a 040-4 scale) unattainable.
@@ 2.1 From Standard LLM Alignment To Target Reaching Formulation
$\vskip-4.73611pt\max_{\theta}\mathbb{E}_{x\sim\pi_{\theta}}[R(x)],$ (1)
@@ 3.1 Value Function Training via Temporal Difference Learning
In this formulation, $s_{t+n}$ denotes the state reached $n$ steps after $s_{t}$ (in our setting, $s_{t}:=h_{t}$), where $T$ represents the total sequence length. The term $V_{\phi}(s_{t+n})$ serves as a bootstrap estimate of the eventual terminal score starting from that future state, while $r^{T}$ is the final, episode-level reward for the completed sequence. The decay factor $\lambda$ trades off short-horizon bootstrapping against reliance on the terminal Monte Carlo target—approaching pure MC as $\lambda\to 1$. The value function is then trained to minimize the mean squared error between its predictions and the generalized returns:
@@ 3.2 Test-time Intervention for Target Attribute Intensity Control
With a trained value function that can predict attribute intensity scores from partial generations, we leverage it to guide the language model toward generating text with a specific attribute score through targeted representation editing. Unlike previous approaches [15, 18, 21, 48] that merely push the model to maximize or minimize a preference, our method enables precise targeting of any scores within the full range of attribute intensities.
@@ 3.3 Efficient Pareto Frontier Approximation
$\vskip-4.30554pt\mathbf{t}=I(\mathbf{n}_{1},\mathbf{n}_{2},\beta),\quad\beta\in[0,1]$ (11)
@@
The core of our approach is using our precise attribute intensity control capability to directly generate samples at specific target points along the Pareto frontier, which traditional methods cannot achieve. For each iteration, we identify the most promising target by calculating the coverage gap at each point $\mathbf{t}$ as:
$\vskip-4.30554ptG(\mathbf{t},\mathcal{N})=\min_{\mathbf{n}\in\mathcal{N}}|\mathbf{t}-\mathbf{n}|_{2}.$ (12)
@@ 4.2 Main Results
We evaluate the effectiveness of Pre-Control for precise attribute intensity control on HelpSteer2 and Code-UltraFeedback. Table 1 presents the main results on both relative positive and negative target vectors, which illustrate Pre-Control’s bidirectional finer-grained control capability. Crucially, the strong performance of our method is not limited to these specific points. We provide a comprehensive evaluation across a wide range of target scores in Appendix B.3, with full results in Table 14 and  15, which confirms the robustness and consistency of our findings. To better demonstrate the effectiveness of our approach, we present a comparison of attribute intensity distributions before and after our intervention in Figure 5.
@@ 5.1 LLM Alignment
Alignment approaches for LLMs fall into two primary paradigms. Fine‐tuning via RLHF [36, 52, 39]—where a reward model guides policy optimization—yields robust performance but depends on a multi‐stage loop of reward learning, policy updates, and rollouts, which can be resource‐intensive [36, 39]. Direct Preference Optimization (DPO) [34] recasts this as a supervised loss, removing the need for online sampling, yet still demands significant memory to maintain both policy and reference models. Inference‐time interventions sidestep model updates: prompt engineering—crafting instructions (with or without examples)—can nudge outputs toward desired behaviors with almost no extra compute [2]. Guided decoding as another effective branch has also been well-explored: ARGS weave reward‐model scores into token probabilities to steer generation [16]. Mudgal et al. [28] and Han et al. [14] train a prefix‑based reward scorer to guide generation from a partial hypothesis. DeAL [15], by contrast, casts decoding as an A* search, using heuristic costs to optimize token selection. Transfer Q [7] introduces an inference‐time policy adjustment. Energy-based methods like COLD  [33] and BOLT  [26] perform iterative gradient-based search at the logit level to find low-energy (high-reward) sequences. However, these methods fundamentally lack a principled mechanism for precise attribute intensity control. They are designed to monotonically shift outputs toward preferred extremes—for example, by maximizing a reward function (i.e. ARGS) or minimizing a static energy function (i.e. COLD and BOLT). They are not formulated as a target-reaching problem and thus offer no built-in stopping criterion to hit a specific scalar target (e.g., “helpfulness = 3“), which would require extensive, per-sample hyperparameter tuning.
@@
Another critical direction in LLM alignment is multi-objective alignment, which is crucial for real-world deployment where LLMs must balance competing attributes based on user preferences. Recent works on multi-objective alignment have explored various ways to approximate Pareto-optimal trade-offs. Rame et al. [35] trains separate policies for each reward preference via RLHF and interpolates them post hoc. MODPO [51] extends Direct Preference Optimization to handle multiple objectives. RiC [47] reduces training costs by applying reward-conditioned supervised fine-tuning and lightweight online data augmentation. Panacea [50] further embeds preference vectors into model parameters through SVD-LoRA, enabling a single model to generalize across objectives after training. CPO [13] introduces controllable preference tokens and extends SFT/DPO to condition explicitly on multi-objective preference scores, enabling controllable trade-offs among values such as helpfulness, honesty, and harmlessness. Preference Orchestrator [22] learns a prompt-aware adapter that infers context-specific preference weights to combine multiple reward models, automatically adapting objective trade-offs to each input. PARM Lin et al. [20] tackles multi-objective test-time alignment by training a single preference-aware autoregressive reward model conditioned on preference vectors to guide frozen LLMs along preference trade-offs during decoding.Despite these advancements, these methods rely on costly retraining to inject various multi-objective preferences. In contrast, our method achieves efficient alignment entirely at test time.
@@ 5.2 Representation Engineering
Activation perturbation began with plug-and-play methods like PPLM  [8], , which use attribute-specific classifiers to nudge hidden states. However, this approach is often computationally inefficient for large LMs, as its intervention loop requires repeated forward and backward passes through the expensive LM head to compute gradients from the logit-space loss  [6, 27]. Subsequent studies showed that both learned and handcrafted steering vectors can control style [37, 40] and that targeting attention‑head outputs boosts factual accuracy [19]. Panickssery et al. [32] applies steering vector, constructed from residual-stream activation differences between positive and negative exemplars, at inference to intervene behaviors. Cao et al. [4] optimizes steering vector using contrastive human-preference pair and use it to inject personalized control without additional model training. Liu et al. [24] further interpret in‑context learning as shifting latent states toward task‑relevant regions. More recently, representation fine‑tuning leverages low‑rank projection matrices to edit activations efficiently, often matching or outperforming parameter‑efficient tuning [44, 45], whereas  [25]’s two‑phase approach first identifies steering directions via fine‑tuning before applying them, adding extra complexity. Similarly, these methods primarily focus on binary or categorical attribute control rather than precisely targeting specific attribute intensities on a continuous scale. Similarly, these methods primarily focus on binary or categorical attribute control. They are not designed for precisely targeting specific attribute intensities on a continuous scale, as they lack the explicit target-reaching objective that provides a principled, per-sample steering signal and stopping criterion.
@@ 7 Reproducibility Statement
We release code, configs, and scripts at https://github.com/Pre-Control/pre-control. The core algorithmic details are specified in Section 3 and Algorithm 1, and the full experimental setup appears in Section 4. Dataset descriptions and preprocessing steps for HelpSteer2 and Code-UltraFeedback are provided in Appendix C.1 and Appendix C.2, respectively. Implementation specifics—including model choices, intervention layer, value-function architecture, and training targets (ArmoRM), hyperparameters, random seeds, and decoding settings—are documented in Appendix C.3 and in the released configuration files. Our Pareto-frontier construction and interpolation procedures, along with the hypervolume, sparsity, and #PF computations, are detailed in Appendix C.4; metric definitions (Self-BLEU, $\ell_{1}$ distance to target, and Success Rate with filtering rules) are summarized in the Metrics subsection of Section 4 and mirrored in the repository’s evaluation scripts. Computing infrastructure (hardware, GPU hours, and environment) is reported in Appendix E.
@@ B.3 Full Intervention Results
To demonstrate the robustness of our approach, we assess its performance across a range of target attribute intensity scores. The complete results for these varied targets are shown in Table 14 and 15. Pre-Control demonstrates a consistent, strong performance compared to all baselines in various settings.
@@ C.4.2 Upper Convex Hull Interpolation
Appendix D Hyperparameter Study
To better understand the characteristics of Pre-Control, we analyze the sensitivity of its key hyperparameter, intervention step size $\alpha$ (in Eq. 9), which controls the magnitude of the gradient update at each intervention step. We vary $\alpha$ across several orders of magnitude and measure the impact on the overall Success Rate. We present the result in Figure 12. Success Rate remains stable, fluctuating between 6.51% and 6.99% for $\alpha$ values ranging from $10^{-1}$ down to $10^{-3}$. This demonstrates that Pre-Control is robust across a wide range of step size values.
Figure 12: Sensitivity of intervention step size $\alpha$.
Appendix E Computing Infrastructure
@@
Appendix F Latency
We analyze the computational efficiency of Pre-Control by evaluating its two primary costs: the one-time, offline training of the value function and online inference-time intervention.
Value Function Training
Training a 4-layer neural network on 1,0162 samples requires only 0.34 GPU hours. This modest, one-time cost makes the value function practical to train and update.
Inference-Time Intervention
To quantify the computational overhead of test-time intervention, we compared the cost of generating 434 test samples from HelpSteer2 using LLaMA-3.2-3b against all baselines. We present the results in Table 10. While Pre-Control incurs moderate overhead due to token-level optimization, it remains within the same order of magnitude as other baselines. Given its strong performance in multi-attribute controllability, this overhead represents a practical trade-off.
Since generation steps (output token length) and the degree of parallelization (batch size) significantly impact practical computational cost, we analyzed how overhead scales with each factor independently. For output token length, we keep batch size the same to 24. For batch size, we keep the output token length the same to 512. Results in Table 10 indicates that the inference cost scales linearly with the output token length while also benefiting significantly from parallelization, as the total cost decreases with a larger batch size.
Method GPU Hours
Base 0.02
Prompting 0.02
ITI 0.06
MAT-Steer 0.07
RE-Control [^5] RE-Control’s computational cost is sensitive to the number of intervention steps. We follow [18] to use 100 steps which yields the best performance in terms of reward scores. 0.10
Pre-Control 0.09
Table 9: Computational Cost Comparison
of LLaMA-3.2-3b on HelpSteer2
Output Token Length GPU Hours
256 0.04
512 0.09
768 0.14
Batch Size GPU Hours
12 0.14
24 0.09
48 0.05
Table 10: Computational Cost on Different
Output Token Length and Batch Size
Appendix G Case Study
In Table˜11, Table˜12, and Table˜13, we present qualitative examples demonstrating Pre-Control’s ability to precisely control attribute intensities.
@@ Appendix G Case Study
Positive Target Scenario 1. The base model generates a response with scores [4,4,4,1,1] for helpfulness, coherence, correctness, complexity, and verbosity, respectively. While the response is helpful and correct, it lacks detail—providing only minimal explanations without elaborating on command options or their purposes. To address this deficiency, we set target scores of [4,4,4,2,2], aiming to maintain the high quality while increasing both complexity and verbosity to provide more comprehensive information. After applying Pre-Control, the model successfully achieves target scores by enriching the response with explicit clarifications of command flags, detailed option descriptions, and expanded explanations of each component’s purpose.
Positive Target Scenario 2. The base model produces a functionally correct solution with scores [3,3,2,3,3] for complexity and efficiency, style, explanation, instruction-following, and readability, but its explanation of the code is brief and offers limited guidance beyond the raw implementation. We set a higher target on the explanation dimension and apply Pre-Control, which yields a response with scores [3,3,3,3,3] by enriching the answer with an explicit step-by-step explanation clarifying the function’s behavior and usage. This example illustrates that Pre-Control can selectively improve explanatory quality while preserving the underlying correctness and overall structure of the original solution.
@@
Positive Target Example 1 Prompt:
@@
Positive Target Example 2 Prompt:
USER: Author a Python procedure labeled ćoncatenateAlphabetsẃhich accepts two parameters of the data type string and outputs the resultant string, amalgamated from both inputs. string1 = "Hello" string2 = "World" ASSISTANT:
Base To create a Python procedure named ‘concatenateAlphabets‘ that takes two string parameters and returns their concatenated result, you can define a function as follows:
    ‘‘‘python
    def concatenateAlphabets(string1, string2):
        Concatenates two input strings and returns the result.
        Parameters:
        string1 (str): The first input string.
        string2 (str): The second input string.
        Returns:
        str: The concatenated result of string1 and string2.
        return string1 + string2
    # Example usage:
    string1 = "Hello"
    string2 = "World"
    result = concatenateAlphabets(string1, string2)
    print(result)  # Output: HelloWorld
This function simply uses the ‘+‘ operator to concatenate the two input strings and returns the combined result. You can test it with the provided example strings "Hello" and "World", which will output "HelloWorld".
Pre-Control To create a Python function named ‘concatenateAlphabets‘ that takes two string parameters and returns their concatenated result, you can follow the code example below. This function will simply use the ‘+‘ operator to concatenate the two input strings. Here’s how you can implement it:
    ‘‘python
    def concatenateAlphabets(string1, string2):
        Concatenates two input strings and returns the result.
        Parameters:
        string1 (str): The first input string.
        string2 (str): The second input string.
        Returns:
        str: The concatenated result of string1 and string2.
        return string1 + string2
    # Example usage:
    string1 = "Hello"
    string2 = "World"
    result = concatenateAlphabets(string1, string2)
    print(result)  # Output: HelloWorld
Explanation: The function ‘concatenateAlphabets‘ is defined with two parameters, ‘string1‘ and ‘string2‘. Inside the function, the ‘+‘ operator is used to concatenate ‘string1‘ and ‘string2‘. The concatenated result is returned. An example usage is provided to demonstrate how the function works with the strings ‘"Hello"‘ and ‘"World"‘, producing the output ‘"HelloWorld"‘.
Table 13: Qualitative examples of positive target score showing the alignment performance of Pre-Control. Base response has a score of 3,3,2,3,3. Pre-Control response has a score of 3,3,3,3,3.